OpenStudy (anonymous):

i don't really understand how to solve this problem can somebody help me pwease Amelia is planning the trajectory of her next flight using the altitude and distance from Paris, France. She has determined her function to be f(x) = -4x + 102. Based on the situation above, describe any necessary restrictions to the domain and range. Provide justification for the restrictions.

OpenStudy (anonymous):

find the x

OpenStudy (anonymous):

ok and how do you do that like can you show me the steps

OpenStudy (anonymous):

Im still not reallyunderstanding it

OpenStudy (anonymous):

wait one min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

ok but i dont think its answering this question Based on the situation above, describe any necessary restrictions to the domain and range. Provide justification for the restrictions

OpenStudy (abdela25):

do you want to graph it or write it as a function(equation) which means solving fro x to complete any restrictions

OpenStudy (anonymous):

write it as a function

OpenStudy (abdela25):

not necessarily a fraction this can kind of be confusing it is better to solve for x not using fractions

OpenStudy (anonymous):

so are we going to use a fraction then?

OpenStudy (abdela25):

no fractions are confusing we will solve for x first then solve the function

OpenStudy (anonymous):

oh ok

OpenStudy (abdela25):

first i will give you the answer and see if you know how i got it and explain it to you

OpenStudy (anonymous):

abdela25

OpenStudy (anonymous):

ok

OpenStudy (abdela25):

the answer is x=51/2 do you know how i got that?

OpenStudy (anonymous):

hold on im tryin to figure it out now

OpenStudy (abdela25):

okay take your time

OpenStudy (anonymous):

how did you get the answer????

OpenStudy (abdela25):

ok first i isolated the terms with 4x to the left hand side by subtracting 102 from each side so it will look like this: -(4x) + (102-102) = -102

OpenStudy (abdela25):

do you understand what i did here

OpenStudy (anonymous):

I dont

OpenStudy (anonymous):

me either

OpenStudy (abdela25):

okay well i had to isolate terms with 4x to the left so i subtracted 102 from each side

OpenStudy (abdela25):

@thatgirl;] @prettigirlrock0804

OpenStudy (anonymous):

ohh ok now i understand

OpenStudy (abdela25):

102-102 = 0 so this is nothing -4x - 102 is not possible becuase 4 has a variable - X so it will look like this: -4x=-102 (102 is negative becuase it went on the other side) do you guys understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@abdela25 I understand that but does it solve the question

OpenStudy (abdela25):

yeah of course it does we have to solve for x first when we have x we replace x in the question by what ever the answer is

OpenStudy (abdela25):

then easy multiplication from there

OpenStudy (anonymous):

so we have right now -4x=-102 right

OpenStudy (abdela25):

yeah

OpenStudy (anonymous):

when you divde you get -25.5

OpenStudy (abdela25):

so we have to put X by itself since it is multiplied with the 4 we have to divide 4 from each side

OpenStudy (anonymous):

and you get -25.5

OpenStudy (abdela25):

no -4/-4 = 1x and -102/-4 is canceled out

OpenStudy (abdela25):

you would get -25.5 if there was only one negative number but both number are negative -102/-4 both negative signs cancel out each other so it is left at 102/4 (both positive now)

OpenStudy (abdela25):

do you understand or should i go back

OpenStudy (anonymous):

ok so you would divde 102/4 right

OpenStudy (abdela25):

not divide but reduce it

OpenStudy (abdela25):

102/4 can be simplified still

OpenStudy (abdela25):

you reduce by finding the GCD of both numbers

OpenStudy (anonymous):

ohh ok so it would be 51/2

OpenStudy (abdela25):

yes thats the answer! thats what i got in the beginning before explaining

OpenStudy (abdela25):

so X=51/2

OpenStudy (anonymous):

oh ok i finally got it thank you so muchh:) do you think you can help me with one more question

OpenStudy (abdela25):

welcomes! and yeah i think i got some more time whats the question?

OpenStudy (anonymous):

Neveah has created the function f(x) = the quantity of 2x plus 4, divided by 3 to represent the growth of her hair, where x represents the number of months since her last hair cut. Neveah discovers that using the inverse function to solve for x = 6, she can predict when her hair will have grown 6 inches. Explain to Neveah how to accomplish this, using complete sentences.

OpenStudy (abdela25):

ok so this problem wants to use words and not an equation right?

OpenStudy (anonymous):

right

OpenStudy (abdela25):

okay something went wrong Either you typed it wrong, or the textbook got it wrong. By solving for x=6, she'll end up figuring out how long her hair will be in six months, not the time her hair reaches six inches. You meant to say, "solve for f(x) = 6." :)

OpenStudy (abdela25):

cuz the way you said typed it there is no possible answer to it it must be f(x) = 6 which is simple algebra and wil be done in 2 minutes or less

OpenStudy (abdela25):

to figure out how long her hair will grow in 6 months rather than 6 in

OpenStudy (abdela25):

you even said it yourself "x represents the number of months since her last hair cut...x = 6" this clearly means 6 months and not 6 in in how ever many months cuz they tell us 6 months already theyre asking fro f(x)=6

OpenStudy (abdela25):

are you with me?

OpenStudy (abdela25):

?????

OpenStudy (abdela25):

if i am correct here is the answer: simple algebra. (2x+4)/3 = 6. Solve for x. That's an Algebra 1 problem, easy. Converting it to complete sentences is left as an exercise to the reader. ;) any confusion ???

OpenStudy (anonymous):

yes im still with you

OpenStudy (abdela25):

okay welll do you get it now

OpenStudy (abdela25):

the question was printed out wrong or typed wrong cuz there was no possible answer to the way you typed it but my way which is removing 6 in and keeping it 6 months gives you the answer of simple algebra. (2x+4)/3 = 6. Solve for x which is very simple Converting it to complete sentences is left as an exercise to the reader

OpenStudy (anonymous):

ok so like do you multiply 6 and 3 and idk about the question i mean that is what its askin me to do

OpenStudy (abdela25):

yeah its a typo small little thing i did the same thing

OpenStudy (abdela25):

(2x+4)/3 = 6 solve for X

OpenStudy (anonymous):

ok so this is what i got 2x+4=18 2x=14 x=7 is that right

OpenStudy (abdela25):

yep x=7 is the correct answer

OpenStudy (anonymous):

ok thank you for all your help most appreciated:)

OpenStudy (abdela25):

your welcome! any more questions?